# KBQA-R1: Reinforcing Large Language Models for Knowledge Base Question Answering

### v1

@@ IV-C2 Policy Optimization (GRPO)
We optimize the policy $\pi_{\theta}$ using Grouped-Reward Policy Optimization (GRPO) [shao2024deepseekmath], a PPO [schulman2017proximal] variant that leverages a low-variance, per-prompt advantage estimation without requiring a learned value function. The overall objective maximizes the expected advantage, regularized by a KL-divergence term against a frozen reference policy $\pi_{\text{ref}}$ to ensure training stability [ouyang2022training]:
@@
The key feature of GRPO is its definition of the advantage function $\hat{A}_{t}$. For a given prompt $x$, we execute $n$ rollouts with the current policy $\pi_{\theta}$ to generate $n$ candidate trajectories $\{y_{i}\}_{i=1}^{n}$ and their corresponding scalar rewards $\{r_{i}\}_{i=1}^{n}$. Instead of using a learned value function (as in standard actor-critic methods [schulman2015high]), GRPO computes the advantage by centering the rewards within this group, using the group’s mean reward as a baseline:
@@
This per-prompt baseline significantly reduces reward variance, a technique that echoes the REINFORCE with baseline approach [REINFORCE]. The final loss function integrates this advantage estimate, $\hat{A}_{i}$, with standard PPO mechanisms like value loss (if used) and clipping for robust optimization. Rollouts are efficiently executed using vLLM [vLLM] with top-p/temperature sampling, and throughput is maximized via dynamic batching.
@@ V-A1 Datasets
We conduct experiments on three widely-used KBQA benchmarks, each designed to evaluate different aspects of model generalization and reasoning capabilities. All datasets are grounded on Freebase [Freebase]. GrailQA [GrailQA] is a large-scale dataset specifically designed to evaluate KBQA models across three generalization levels: i.i.d., compositional, and zero-shot. It contains 64,331 questions in total, with 44,337 training questions, 13,231 validation questions and 6,763 test questions. Following prior work [ChatKBQA, luo2025kbqa], we use the dev set for evaluation. The compositional and zero-shot settings are particularly challenging, requiring models to handle unseen combinations of entities and relations. WebQSP [WebQSP] is an enriched version of WebQuestions, providing semantic parses for 4,737 questions. The dataset is split into 3,098 training questions and 1,639 test questions. GraphQuestions [GraphQ] tests KBQA models on complex graph-structured reasoning. It contains 5,166 questions in total, with 2,508 for training and 2,658 for testing. The dataset challenges models to navigate multi-hop relationships.
@@ V-A2 Baselines
Finetune-based Methods. These methods are trained on the complete training datasets and serve as upper-bound references: RnG-KBQA [RnG-KBQA] employs a retrieve-and-generate framework that first retrieves relevant knowledge from the KB and then generates executable logical forms. DecAF [DecAF] uses multi-granular retrieval strategies to ensure robust KBQA performance by progressively refining retrieved knowledge. TIARA [TIARA] is a semantic-parsing-based approach that maps questions to structured queries through iterative refinement. KBQA-o1 [luo2025kbqa] is a recent MCTS-based agentic KBQA method that employs heuristic search with policy and reward models.
For GraphQuestions, following KBQA-o1’s setup, we also compare with SPARQA [SPARQA], BERT+Ranking [GrailQA], and ArcaneQA [ArcaneQA].
Prompt Based Methods. These methods operate under the same limited annotation constraint as KBQA-R1: KB-BINDER [KB-BINDER] leverages in-context learning with GPT-3.5-turbo to bind questions to KB entities and relations. KB-Coder [KB-Coder] adopts a code-style in-context learning approach to generate logical forms with GPT-3.5-turbo. ARG-KBQA [ARG-KBQA] uses augmented reasoning graphs with GPT-3.5-turbo for improved question answering.
@@ V-A3 Evaluation Metrics
KB-BINDER [KB-BINDER] Codex-davinci-002 40.0 43.3 33.9 36.6 40.1 44.0 38.7 42.2
KB-Coder [KB-Coder] GPT-3.5-turbo 40.6 45.5 34.5 38.6 42.2 47.3 40.1 44.9
ARG-KBQA [ARG-KBQA] GPT-3.5-turbo 46.6 51.5 36.4 41.8 46.6 52.1 43.8 48.5
@@
RnG-KBQA [RnG-KBQA] T5-large 86.7 89.0 61.7 68.9 68.8 74.7 69.5 76.9
DecAF [DecAF] T5-large 88.7 92.4 71.5 79.8 65.9 77.3 72.5 81.4
TIARA [TIARA] T5-large 88.4 91.2 66.4 74.8 73.3 80.7 75.3 81.9
KBQA-o1 [luo2025kbqa] Llama-3.1-8B 77.8 $\pm$0.5 85.5 $\pm$0.4 76.3 $\pm$0.6 77.6 $\pm$0.5 68.1 $\pm$0.8 76.1 $\pm$0.4 71.9 $\pm$0.3 78.5 $\pm$1.0
@@
KB-BINDER [KB-BINDER] Codex-davinci-002 52.6
KB-Coder [KB-Coder] GPT-3.5-turbo 55.7
ARG-KBQA [ARG-KBQA] GPT-3.5-turbo 58.8
Interactive-KBQA [Interactive-KBQA] GPT-4-turbo 71.2
@@
RnG-KBQA [RnG-KBQA] T5-large 75.6
DecAF [DecAF] T5-large 76.7
TIARA [TIARA] T5-large 78.9
MCTS-KBQA [MCTS-KBQA] Llama-3.1-8B 76.0
KBQA-o1 [luo2025kbqa] Llama-3.1-8B 57.8
@@
KB-BINDER [KB-BINDER] Codex-davinci-002 27.1
KB-Coder [KB-Coder] GPT-3.5-turbo 31.1
@@
SPARQA [SPARQA] BERT-base 21.5
BERT+Ranking [GrailQA] BERT-base 25.0
ArcaneQA [ArcaneQA] BERT-base 31.8
CoTKR [CoTKR] Llama-3-8B 47.5
KBQA-o1 [luo2025kbqa] Llama-3.1-8B 48.7
@@ V-A4 Training Setup
Model Architecture. We use Llama-3.1-8B-Instruct as the default backbone. For fair comparison with KBQA-o1 [luo2025kbqa], all experiments use the same base model architecture.
@@
GRPO Configuration. We adopt the GRPO algorithm [shao2024deepseekmath] with the following hyperparameters: (1) Rollout sampling: $n=5$ responses per prompt with temperature $\tau=1.0$ and top-$p=0.99$; (2) Clipping: asymmetric clip ratios $\epsilon_{\text{low}}=0.2$, $\epsilon_{\text{high}}=0.28$ following DAPO [yu2025dapo]; (3) KL regularization: KL loss coefficient $\beta=0.001$; (4) Reward weights and gating: we set $\lambda_{\text{outcome}}=1.0$ and $\lambda_{\text{format}}=0.1$, and use RRCG confidence thresholds $\tau_{\text{high}}=0.95$ and $\tau_{\text{low}}=0.3$ across all datasets; (5) Batch configuration: train batch size 256, PPO mini-batch size 128, with dynamic micro-batching enabled.
@@ V-C Ablation Study
Following the experimental setup of KBQA-o1 [luo2025kbqa], we conduct a focused comparison among methods that share the same Llama-3.1-8B backbone and Freebase execution environment. The compared baselines can be grouped into three categories. (1) End-to-end generation methods: RG-E2E and GR-E2E are adapted from DecAF [DecAF] and ChatKBQA [ChatKBQA], respectively. RG-E2E follows a retrieve-then-generate paradigm, while GR-E2E first generates a preliminary logical form and then refines it with KB retrieval. (2) Step-by-step prompting methods: CoT-SbS and ToT-SbS are implemented by instantiating the CoT-based QueryAgent [QueryAgent] and the ToT-based ToG framework [ToG] on Llama-3.1-8B, prompting the model to alternate between intermediate thoughts and KB queries. (3) MCTS-based agentic method: MCTS corresponds to the MCTS-optimized variant in KBQA-o1 [luo2025kbqa] without incremental Finetuning. Figure 3(a) visualizes F1 scores across six evaluation dimensions. KBQA-R1 achieves the largest coverage area, demonstrating superior overall performance across all settings, with the most pronounced gap in zero-shot dimensions. This validates our hypothesis that RL-based training fosters more robust reasoning capabilities than SFT. In contrast, end-to-end and step-by-step baselines cluster in the inner region, reflecting limited generalization. Figure 3(b) further breaks down performance by logical operation type. KBQA-R1 dominates across all categories, showing particularly significant advantages in complex operations. Conversely, baselines struggle with rare operations, underscoring their inability to generalize to infrequent query patterns.
@@ VI Conclusion
VII Biography Section
[Uncaptioned image] Xin Sun is a joint Ph.D. candidate from University of Science and Technology of China(USTC) and Institute of Automation, Chinese Academy of Sciences(CASIA). He received his bachelor degree from Shanghai Jiao Tong University(SJTU). His current research interests mainly include trustworthy learning and information retrieval. He has published papers in top-tier conferences as first author, such as ACL, KDD, EMNLP, etc.
[Uncaptioned image] Zhongqi Chen received the master’s degree in Control Science and Engineering from Xi’an Jiaotong University in 2022. He joined Ant Group the same year, where he is currently a Senior Algorithm Engineer in the Anti-Money-Laundering Algorithm Team. His research focuses on deep learning, text generation, and post-training of large language models, with applications in financial security and intelligent risk detection.
[Uncaptioned image] Xing Zheng received the B.Eng. and M.Eng. degrees in Information and Communication Engineering from Huazhong University of Science and Technology in 2017 and 2020, respectively. Since 2020, he has been with Ant Group as a Senior Algorithm Engineer, working on large language model data science and financial security risk control. His research focuses on natural language processing, data evaluation for large language models, and post-training optimization.
[Uncaptioned image] Qiang Liu (Member, IEEE) is an Associate Professor with the New Laboratory of Pattern Recognition (NLPR), State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS), Institute of Automation, Chinese Academy of Sciences (CASIA). He received his PhD degree from CASIA. Currently, his research interests include data mining, misinformation detection, LLM safety and AI for science. He has published papers in top-tier journals and conferences, such as IEEE TKDE, AAAI, NeurIPS, KDD, WWW, SIGIR, CIKM, ICDM, ACL and EMNLP.
[Uncaptioned image] Shu Wu (Senior Member, IEEE) received his B.S. degree from Hunan University, China, in 2004, M.S. degree from Xiamen University, China, in 2007, and Ph.D. degree from Department of Computer Science, University of Sherbrooke, Quebec, Canada, all in computer science. He is a Professor with the New Laboratory of Pattern Recognition (NLPR), State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS), Institute of Automation, Chinese Academy of Sciences (CASIA). He has published more than 50 papers in the areas of data mining and information retrieval in international journals and conferences, such as IEEE TKDE, IEEE THMS, AAAI, ICDM, SIGIR, and CIKM. His research interests include data mining, information retrieval, and recommendation.
[Uncaptioned image] Bowen Song received the Ph.D. degree in applied math and statistics from Stony Brook University in 2015. He joined Ant Group in 2017, where he is currently a Senior Staff Algorithm Engineer with the Anti-Money-Laundering Algorithm Team. His research interests include behavior sequential learning, deep graph learning, and their applications in financial risk management and web3.
[Uncaptioned image] Weiqiang Wang (Member, IEEE) received the Ph.D. degree from the University of Southern California, USA, in 2008. From 2007 to 2010, he was a Research Fellow at the University of Southern California. He is currently working as the Team Leader of the Tiansuan Lab, Ant Group, China. He has published more than 90 articles in top-tier international conferences and journals. His research interests include graph learning, computer vision, and natural language processing.
[Uncaptioned image] Zilei Wang (Member, IEEE) received the BS and PhD degrees in control science and engineering from the University of Science and Technology of China (USTC), Hefei, China, in 2002 and 2007, respectively. He is currently an Associate Professor with the Department of Automation, USTC, where he is the Founding Leader of the Vision and Multimedia Research Group. His research interests include computer vision, multimedia, and deep learning. Dr. Wang is a member of the Youth Innovation Promotion Association, Chinese Academy of Sciences.
[Uncaptioned image] Liang Wang (Fellow, IEEE) received both the BEng and MEng degrees from Anhui University in 1997 and 2000, respectively, and the PhD degree from the Institute of Automation, Chinese Academy of Sciences (CASIA) in 2004. Currently, he is a full professor of the Hundred Talents Program at the State Key Laboratory of Multimodal Artificial Intelligence Systems, CASIA. His major research interests include machine learning, pattern recognition, and computer vision. He has widely published in highly ranked international journals such as IEEE TPAMI and IEEE TIP, and leading international conferences such as CVPR, ICCV, and ECCV. He has served as an Associate Editor of IEEE TPAMI, IEEE TIP, and PR. He is an IEEE Fellow and an IAPR Fellow.

### v2

@@ IV-C2 Policy Optimization (GRPO)
We optimize the policy $\pi_{\theta}$ using Grouped-Reward Policy Optimization (GRPO) [32], a PPO [31] variant that leverages a low-variance, per-prompt advantage estimation without requiring a learned value function. The overall objective maximizes the expected advantage, regularized by a KL-divergence term against a frozen reference policy $\pi_{\text{ref}}$ to ensure training stability [26]:
@@
The key feature of GRPO is its definition of the advantage function $\hat{A}_{t}$. For a given prompt $x$, we execute $n$ rollouts with the current policy $\pi_{\theta}$ to generate $n$ candidate trajectories $\{y_{i}\}_{i=1}^{n}$ and their corresponding scalar rewards $\{r_{i}\}_{i=1}^{n}$. Instead of using a learned value function (as in standard actor-critic methods [30]), GRPO computes the advantage by centering the rewards within this group, using the group’s mean reward as a baseline:
@@
This per-prompt baseline significantly reduces reward variance, a technique that echoes the REINFORCE with baseline approach [42]. The final loss function integrates this advantage estimate, $\hat{A}_{i}$, with standard PPO mechanisms like value loss (if used) and clipping for robust optimization. Rollouts are efficiently executed using vLLM [17] with top-p/temperature sampling, and throughput is maximized via dynamic batching.
@@ V-A1 Datasets
We conduct experiments on three widely-used KBQA benchmarks, each designed to evaluate different aspects of model generalization and reasoning capabilities. All datasets are grounded on Freebase [1]. GrailQA [5] is a large-scale dataset specifically designed to evaluate KBQA models across three generalization levels: i.i.d., compositional, and zero-shot. It contains 64,331 questions in total, with 44,337 training questions, 13,231 validation questions and 6,763 test questions. Following prior work [21, 22], we use the dev set for evaluation. The compositional and zero-shot settings are particularly challenging, requiring models to handle unseen combinations of entities and relations. WebQSP [48] is an enriched version of WebQuestions, providing semantic parses for 4,737 questions. The dataset is split into 3,098 training questions and 1,639 test questions. GraphQuestions [34] tests KBQA models on complex graph-structured reasoning. It contains 5,166 questions in total, with 2,508 for training and 2,658 for testing. The dataset challenges models to navigate multi-hop relationships.
@@ V-A2 Baselines
Finetune-based Methods. These methods are trained on the complete training datasets and serve as upper-bound references: RnG-KBQA [47] employs a retrieve-and-generate framework that first retrieves relevant knowledge from the KB and then generates executable logical forms. DecAF [50] uses multi-granular retrieval strategies to ensure robust KBQA performance by progressively refining retrieved knowledge. TIARA [33] is a semantic-parsing-based approach that maps questions to structured queries through iterative refinement. KBQA-o1 [22] is a recent MCTS-based agentic KBQA method that employs heuristic search with policy and reward models.
For GraphQuestions, following KBQA-o1’s setup, we also compare with SPARQA [38], BERT+Ranking [5], and ArcaneQA [6].
Prompt Based Methods. These methods operate under the same limited annotation constraint as KBQA-R1: KB-BINDER [19] leverages in-context learning with GPT-3.5-turbo to bind questions to KB entities and relations. KB-Coder [25] adopts a code-style in-context learning approach to generate logical forms with GPT-3.5-turbo. ARG-KBQA [39] uses augmented reasoning graphs with GPT-3.5-turbo for improved question answering.
@@ V-A3 Evaluation Metrics
KB-BINDER [19] Codex-davinci-002 40.0 43.3 33.9 36.6 40.1 44.0 38.7 42.2
KB-Coder [25] GPT-3.5-turbo 40.6 45.5 34.5 38.6 42.2 47.3 40.1 44.9
ARG-KBQA [39] GPT-3.5-turbo 46.6 51.5 36.4 41.8 46.6 52.1 43.8 48.5
@@
RnG-KBQA [47] T5-large 86.7 89.0 61.7 68.9 68.8 74.7 69.5 76.9
DecAF [50] T5-large 88.7 92.4 71.5 79.8 65.9 77.3 72.5 81.4
TIARA [33] T5-large 88.4 91.2 66.4 74.8 73.3 80.7 75.3 81.9
KBQA-o1 [22] Llama-3.1-8B 77.8 $\pm$0.5 85.5 $\pm$0.4 76.3 $\pm$0.6 77.6 $\pm$0.5 68.1 $\pm$0.8 76.1 $\pm$0.4 71.9 $\pm$0.3 78.5 $\pm$1.0
@@
KB-BINDER [19] Codex-davinci-002 52.6
KB-Coder [25] GPT-3.5-turbo 55.7
ARG-KBQA [39] GPT-3.5-turbo 58.8
Interactive-KBQA [43] GPT-4-turbo 71.2
@@
RnG-KBQA [47] T5-large 75.6
DecAF [50] T5-large 76.7
TIARA [33] T5-large 78.9
MCTS-KBQA [44] Llama-3.1-8B 76.0
KBQA-o1 [22] Llama-3.1-8B 57.8
@@
KB-BINDER [19] Codex-davinci-002 27.1
KB-Coder [25] GPT-3.5-turbo 31.1
@@
SPARQA [38] BERT-base 21.5
BERT+Ranking [5] BERT-base 25.0
ArcaneQA [6] BERT-base 31.8
CoTKR [49] Llama-3-8B 47.5
KBQA-o1 [22] Llama-3.1-8B 48.7
@@ V-A4 Training Setup
Model Architecture. We use Llama-3.1-8B-Instruct as the default backbone. For fair comparison with KBQA-o1 [22], all experiments use the same base model architecture.
@@
GRPO Configuration. We adopt the GRPO algorithm [32] with the following hyperparameters: (1) Rollout sampling: $n=5$ responses per prompt with temperature $\tau=1.0$ and top-$p=0.99$; (2) Clipping: asymmetric clip ratios $\epsilon_{\text{low}}=0.2$, $\epsilon_{\text{high}}=0.28$ following DAPO [51]; (3) KL regularization: KL loss coefficient $\beta=0.001$; (4) Reward weights and gating: we set $\lambda_{\text{outcome}}=1.0$ and $\lambda_{\text{format}}=0.1$, and use RRCG confidence thresholds $\tau_{\text{high}}=0.95$ and $\tau_{\text{low}}=0.3$ across all datasets; (5) Batch configuration: train batch size 256, PPO mini-batch size 128, with dynamic micro-batching enabled.
@@ V-C Ablation Study
Following the experimental setup of KBQA-o1 [22], we conduct a focused comparison among methods that share the same Llama-3.1-8B backbone and Freebase execution environment. The compared baselines can be grouped into three categories. (1) End-to-end generation methods: RG-E2E and GR-E2E are adapted from DecAF [50] and ChatKBQA [21], respectively. RG-E2E follows a retrieve-then-generate paradigm, while GR-E2E first generates a preliminary logical form and then refines it with KB retrieval. (2) Step-by-step prompting methods: CoT-SbS and ToT-SbS are implemented by instantiating the CoT-based QueryAgent [10] and the ToT-based ToG framework [37] on Llama-3.1-8B, prompting the model to alternate between intermediate thoughts and KB queries. (3) MCTS-based agentic method: MCTS corresponds to the MCTS-optimized variant in KBQA-o1 [22] without incremental Finetuning. Figure 3(a) visualizes F1 scores across six evaluation dimensions. KBQA-R1 achieves the largest coverage area, demonstrating superior overall performance across all settings, with the most pronounced gap in zero-shot dimensions. This validates our hypothesis that RL-based training fosters more robust reasoning capabilities than SFT. In contrast, end-to-end and step-by-step baselines cluster in the inner region, reflecting limited generalization. Figure 3(b) further breaks down performance by logical operation type. KBQA-R1 dominates across all categories, showing particularly significant advantages in complex operations. Conversely, baselines struggle with rare operations, underscoring their inability to generalize to infrequent query patterns.